OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

A parallel line to y=7x+2 has the same gradient.

OpenStudy (anonymous):

So the gradient would be...?

OpenStudy (nincompoop):

your key here is to know what kind of a line is parallel to another line. one thing you will notice that parallel lines have the same slope so in the problem given, the parallel line that passes through (1, 3) has the same slope as y=7x+2 use that slope and use the point-slope form and plug in the point (1, 3 ) to solve for the equation

OpenStudy (anonymous):

If we want to have two parallel line should know their slope should be equal ! Means : m1=m2 7=m2 so our m is 7 . then if want this line passes through a point use this : y2-y=m(x2-x) x2=1 y2=-3 m=7 have :-3-y=7*1-7x -3-y=7-7x -y=-7x+10 y=7x-10 Understand ?!:D

OpenStudy (anonymous):

@AnnaPopova